# Multi-Source Position and Direction-of-Arrival Estimation Based on Euclidean Distance Matrices

### v1

@@ VI-D Experiment 1: Position Estimation
TABLE II: Average run times of the EDM-based and SRP-based source position and DOA estimation methods.
@@
Two-source DOA estimation 1.1 27.2
In conclusion, the consistently small median errors and small error distributions demonstrate the versatility of the proposed EDM-based multi-source position estimation method across a wide range of configurations and with low computational cost.
VI-E Experiment 2: DOA Estimation
This section compares the DOA estimation accuracy of the proposed EDM-based method and the baseline SRP-based method for the acoustic scenarios with compact microphone arrays described in Section VI-A. To evaluate the performance, we consider the DOA estimation error for each source, defined as
@@ VI-E Experiment 2: DOA Estimation
where the estimated DOA vectors $\hat{\mathbf{v}}_{1}$ or $\hat{\mathbf{v}}_{2}$ are assigned to true DOA vectors ${\mathbf{v}}_{1}$ or ${\mathbf{v}}_{2}$ using greedy assignment. Fig. 6 presents box plots of the DOA estimation errors $\varepsilon_{1}^{\textrm{DOA}}$ and $\varepsilon_{2}^{\textrm{DOA}}$ over 100 simulated scenarios, considering different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Figure 6: Box plots of DOA estimation errors for both sources. The DOA estimation errors are shown for the EDM-based and SRP-based methods, for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid ($d_{\textrm{c}2}=2$ m, 10 cm cube). The red numbers at the top denote the number of results outside of the plotted range.
@@
$d_{\textrm{c}1}[m]$ EDM, $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ EDM, $\overline{\varepsilon}_{2}^{\textrm{DOA}}$ SRP, $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ SRP, $\overline{\varepsilon}_{2}^{\textrm{DOA}}$
$0.5$ 1.2 2.9 1.3 51.1
11 0.8 1.8 1.2 22.8
22 1.1 1.0 2.4 2.5
33 2.3 0.9 4.5 1.8
44 3.6 0.9 11.7 1.7
@@
As can be observed from Table III for all source distances $d_{\textrm{c}1}$, the proposed EDM-based method yields median DOA estimation errors below $4^{\circ}$ for both sources, which are consistently smaller than the median DOA estimation errors obtained by the SRP-PHAT method. In addition, it can be observed from Fig. 6 that the spread of the box plots is smaller for the EDM-based method than for the SRP-based method for all source configurations. It is also interesting to note that for both methods the median DOA estimation error is larger for the source that is farther away from the array centroid. This can be explained by the fact that the amplitudes of the peaks in the GCC-PHAT and SRP-PHAT functionals corresponding to both sources depend on their signal power ratio. This effect is much more pronounced for the SRP-based method than for the EDM-based method. For instance, when source 1 is close to the microphone array ($d_{\textrm{c}1}=0.5$ m), the median DOA estimation error for source 1 is smaller than the median DOA estimation error for source 2 ($1.2^{\circ}$ and $2.9^{\circ}$ for the EDM-based method; $1.3^{\circ}$ and $51.1^{\circ}$ for the SRP-based method). In contrast, when source 1 is far from the microphone array ($d_{\textrm{c}1}=4$ m), the median DOA estimation error for source 1 is larger than the median DOA estimation error for source 2 ($3.6^{\circ}$ and $0.9^{\circ}$ for the EDM-based method; $11.7^{\circ}$ and $1.7^{\circ}$ for the SRP-based method). When both sources are equi-distant ($d_{\textrm{c}1}=d_{\textrm{c}2}=2$ m), both methods achieve similar median DOA estimation errors for both sources ($1.1^{\circ}$ and $1.0^{\circ}$ for the EDM-based method; $2.4^{\circ}$ and $2.5^{\circ}$ for the SRP-based method). For $d_{\textrm{c}1}=0.5$ m, it can be seen that the median DOA estimation errors are also slightly larger than for $d_{\textrm{c}1}=1$ m, reflecting that the far field model assumption becomes less valid for small distances between the source and the microphones.
The average run times of the EDM-based and SRP-based DOA estimation are shown in Table II. The EDM-based DOA method is about 25 times faster than the SRP-based method, which can again be attributed to the large difference in search-space between both methods. Rather than jointly optimizing a functional in two angle variables using all microphone pairs, the EDM-based method completely eliminates the need for continuous variable optimization, using a cost function based on estimated TDOAs between the reference microphone and the other microphones.
@@
Similarly as for the position estimation results, the consistently small median DOA estimation errors and small error distributions demonstrate the versatility of the proposed EDM-based multi-source DOA estimation method across a wide range of source configurations, and with low computational cost.
@@ VII Conclusions
In this paper, we have proposed novel TDOA-based multi-source position and DOA estimation methods that exploit properties of Euclidean distance matrices. For 3D position estimation, the proposed method requires optimizing only a single continuous variable instead of jointly optimizing three position variables as in SRP-based methods. This variable, representing the distance between each source and a reference microphone, is optimized for each combination of candidate TDOA estimates with a cost function based on the eigenvalues of the Gram matrix associated with the EDM. The estimated relative source positions, obtained from the Gram matrices corresponding to the optimal cost function values, are then mapped to absolute source positions by solving an orthogonal Procrustes problem for each source. For 3D DOA estimation, we define a relative coordinate system for each source, with one axis aligned to the DOA vector. The optimal set of candidate TDOA estimates is determined by minimizing a cost function based on the eigenvalues of a rank-reduced Gram matrix, requiring no continuous variable optimization. The source DOA vectors are then estimated by mapping the relative microphone positions, obtained from the rank-reduced Gram matrices, to the absolute microphone positions, for each source. For two sources in a noisy and reverberant environment, experimental results across a wide range of source and microphone configurations, including compact as well as distributed microphone arrays, show that the proposed EDM-based methods consistently outperform the SRP-based methods in terms of position and DOA estimation accuracy. Furthermore, due to the reduction in continuous variables to be optimized, the computational cost of the EDM-based methods is considerably lower than the SRP-based methods.

### v2

@@ VI-D Computational Complexity Analysis
TABLE I: Average run times of the EDM-based and SRP-based source position and DOA estimation methods.
@@
Two-source DOA estimation 1.2 27.2
VI-E Experiment 1: Position Estimation
@@
Figure 3: Box plots of position estimation errors for both sources. The position estimation errors are shown for the EDM-based method (for $C\in\{2,3\}$) and SRP-based method, for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid ($d_{\textrm{c}2}=2$ m, 2 m cube). The red numbers at the top denote the number of results outside of the plotted range.
Table II summarizes the median position estimation errors for both sources.
TABLE II: Median position estimation errors for the EDM-based and SRP-based methods for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Median position estimation error [cm]
EDM ($C=2$) EDM ($C=3$) SRP
$d_{\textrm{c}1}[m]$ $\overline{\varepsilon}_{1}^{\textrm{pos}}$ $\overline{\varepsilon}_{2}^{\textrm{pos}}$ $\overline{\varepsilon}_{1}^{\textrm{pos}}$ $\overline{\varepsilon}_{2}^{\textrm{pos}}$ $\overline{\varepsilon}_{1}^{\textrm{pos}}$ $\overline{\varepsilon}_{2}^{\textrm{pos}}$
0 0.1 0.8 0.1 0.8 58.0 197.2
11 0.2 0.7 0.2 0.7 4.1 136.1
22 0.6 0.7 0.6 0.7 6.4 7.7
33 2.6 0.6 2.3 0.6 200.0 3.7
44 196.1 0.6 9.4 0.6 350.6 3.1
As can be observed from Table II, for all source distances $d_{\textrm{c}1}$ the proposed EDM-based method yields considerably lower median position estimation errors for both sources than the SRP-based method. In addition, it can be observed from Fig. 3 that the spread of the box plots is much smaller for the EDM-based method (for both $C=2$ and $C=3$) than for the SRP-based method for all source configurations. Furthermore, it is interesting to note that both methods behave quite differently for different source configurations. As can be observed from Table II, the median position estimation error for the EDM-based method mainly depends on source distance, i.e., increases for source 1 with increasing $d_{\textrm{c}1}$, whereas it remains relatively constant for source 2 (at $d_{\textrm{c}2}=2$ m). On the other hand, the median position estimation error for the SRP-based method depends both on the absolute distances between the sources and the array centroid as well as on the relative distance between the sources. For instance, it can be clearly observed that the median position estimation error is larger for the source that is farther away from the array centroid (e.g., for $d_{\textrm{c}1}=3$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}(\mathbf{s})=200$ cm and $\overline{\varepsilon}_{2}^{\textrm{pos}}(\mathbf{s})=3.7$ cm). This can be explained by the peaks of the farther source in the SRP functional being overpowered by the peaks of the closer source. In addition, when both sources are close to the microphones, the median position estimation error for both sources is large (e.g., for $d_{\textrm{c}1}=0$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}(\mathbf{s})=58$ cm and $\overline{\varepsilon}_{2}^{\textrm{pos}}(\mathbf{s})=197.2$ cm). This corresponds to [19, 28], where it was shown that the accuracy of SRP-based position estimation degrades for sources located close to the microphone array because the peaks in the SRP-PHAT functional become narrow. This would necessitate a high 3D grid resolution to detect these peaks, which is computationally impractical in most applications. In contrast, the EDM-based method only requires a one-dimensional grid search on the distance variable $\alpha$, which is computationally more efficient to optimize at a high resolution. Comparing the EDM-based methods for different numbers of candidate TDOA estimates, it can be observed that considering $C=3$ candidate TDOA estimates can slightly improve the position estimation accuracy compared to $C=2$, especially when one of the sources is far from the microphones (e.g., for $d_{\textrm{c}1}=4$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}=196.1$ cm for $C=2$ vs $\overline{\varepsilon}_{1}^{\textrm{pos}}=9.4$ cm for $C=3$).
In conclusion, the consistently small median errors and small error distributions demonstrate the versatility of the proposed EDM-based multi-source position estimation method across a wide range of configurations.
VI-F Experiment 2: DOA Estimation
[…]
Figure 4: Box plots of DOA estimation errors for both sources. The DOA estimation errors are shown for the EDM-based method (for $C\in\{2,3\}$) and SRP-based method, for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid ($d_{\textrm{c}2}=2$ m, 10 cm cube). The red numbers at the top denote the number of results outside of the plotted range.
@@ VI-F Experiment 2: DOA Estimation
EDM ($C=2$) EDM ($C=3$) SRP
$d_{\textrm{c}1}[m]$ $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ $\overline{\varepsilon}_{2}^{\textrm{DOA}}$ $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ $\overline{\varepsilon}_{2}^{\textrm{DOA}}$ $\overline{\varepsilon}_{1}^{\textrm{DOA}}$ $\overline{\varepsilon}_{2}^{\textrm{DOA}}$
$0.5$ 1.2 2.9 1.2 3.2 1.3 51.1
11 0.8 1.8 0.8 1.9 1.2 22.8
22 1.1 1.0 1.1 1.0 2.4 2.5
33 2.3 0.9 2.2 0.9 4.5 1.8
44 3.6 0.9 3.7 1.0 11.7 1.7
@@
As can be observed from Table III for all source distances $d_{\textrm{c}1}$, the proposed EDM-based method (for both $C=2$ and $C=3$) yields median DOA estimation errors below $4^{\circ}$ for both sources, which are consistently smaller than the median DOA estimation errors obtained by the SRP-PHAT method. In addition, it can be observed from Fig. 4 that the spread of the box plots is smaller for the EDM-based method than for the SRP-based method for all source configurations. It is also interesting to note that for both methods the median DOA estimation error is larger for the source that is farther away from the array centroid. This can be explained by the fact that the amplitudes of the peaks in the GCC-PHAT and SRP-PHAT functionals corresponding to both sources depend on their signal power ratio. This effect is much more pronounced for the SRP-based method than for the EDM-based method. For instance, when source 1 is close to the microphone array ($d_{\textrm{c}1}=0.5$ m), the median DOA estimation error for source 1 is smaller than the median DOA estimation error for source 2 ($1.2^{\circ}$ and $2.9^{\circ}$ for the EDM-based method ($C=2$); $1.3^{\circ}$ and $51.1^{\circ}$ for the SRP-based method). In contrast, when source 1 is far from the microphone array ($d_{\textrm{c}1}=4$ m), the median DOA estimation error for source 1 is larger than the median DOA estimation error for source 2 ($3.6^{\circ}$ and $0.9^{\circ}$ for the EDM-based method ($C=2$); $11.7^{\circ}$ and $1.7^{\circ}$ for the SRP-based method). When both sources are equi-distant ($d_{\textrm{c}1}=d_{\textrm{c}2}=2$ m), both methods achieve similar median DOA estimation errors for both sources ($1.1^{\circ}$ and $1.0^{\circ}$ for the EDM-based method ($C=2$); $2.4^{\circ}$ and $2.5^{\circ}$ for the SRP-based method). For $d_{\textrm{c}1}=0.5$ m, it can be seen that the median DOA estimation errors are also slightly larger than for $d_{\textrm{c}1}=1$ m, reflecting that the far field model assumption becomes less valid for small distances between the source and the microphones. Comparing the EDM-based methods for different numbers of candidate TDOA estimates, it can be observed that, unlike for spatially distributed microphones, considering $C=3$ candidate TDOA estimates does not improve DOA estimation accuracy compared to $C=2$. This is consistent with the fact that the number of spurious peaks in the GCC-PHAT functions is typically low for compact microphone arrays when only plausible time lags are considered.
@@
Similarly as for the position estimation results, the consistently small median DOA estimation errors and small error distributions demonstrate the versatility of the proposed EDM-based multi-source DOA estimation method across a wide range of source configurations.
@@ VII Conclusions
In this paper, we have proposed novel TDOA-based multi-source position and DOA estimation methods that exploit properties of Euclidean distance matrices. For 3D position estimation, the proposed method requires optimizing only a single continuous variable instead of jointly optimizing three position variables as in SRP-based methods. This variable, representing the distance between each source and a reference microphone, is optimized for each combination of candidate TDOA estimates with a cost function based on the eigenvalues of the Gram matrix associated with the EDM. The estimated relative source positions, obtained from the Gram matrices corresponding to the optimal cost function values, are then mapped to absolute source positions by solving an orthogonal Procrustes problem for each source. For 3D DOA estimation, we define a relative coordinate system for each source, with one axis aligned to the DOA vector. The optimal set of candidate TDOA estimates is determined by minimizing a cost function based on the eigenvalues of a rank-reduced Gram matrix, requiring no continuous variable optimization. The source DOA vectors are then estimated by mapping the relative microphone positions, obtained from the rank-reduced Gram matrices, to the absolute microphone positions, for each source. For two sources in a noisy and reverberant environment, experimental results across a wide range of source and microphone configurations, including compact as well as distributed microphone arrays, show that the proposed EDM-based methods consistently outperform the SRP-based methods in terms of position and DOA estimation accuracy. Although the computational complexities of the EDM-based methods grow combinatorially with the number of microphones, due to the reduction of the number of continuous variables to be jointly optimized, the run times of the EDM-based methods are lower than the SRP-based methods for the considered experimental setups with six microphones. In future work, it would be interesting to combine the estimated 3D positions or DOAs from the EDM-based methods with source tracking methods [48, 18, 16] for moving source scenarios.